OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to the nearest tenth, if necessary. x2 + 5x - 84 = 0

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Ive tried that and i end up with 20. 135.... but thats not right

OpenStudy (unklerhaukus):

\[x=\frac{-5\pm\sqrt{5^2-4\times 1\times-84}}{2}\] \[=\frac{-5\pm\sqrt{25+336}}{2}\]\[=\frac{-5\pm\sqrt{361}}{2}\]\[=\frac{-5\pm19}{2}\]\[x=7,-12\]